OpenStudy (anonymous):

# Alex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 5 cm and a height of 11 cm. The dimensions of the actual water tanks that will be installed in the park are 9 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm2. Use 22 over 7 for pi. Do not type a comma in your response. For example, an answer of 1,000 would be given as 1000. Answer

OpenStudy (anas.p):

You will need a paper and pencil for this. We will compare the ratio of the two figures. (Model and Original) We will take Model/Original So it will become 5/5*9 5/45 Let this ratio be K We want a ratio for the cross sectional area, right? And cross sectional area is in \[cm ^{2}\] We have K in cm because we compared cm with cm So to have K in cm^2 we take\[K ^{2}\] as the ratio. Now we take out the surface area of the model 2 pi r^2 + 2 pi r h \[(2\times22\div7\times5^{2})+(2\times22\div7\times 5 \times 11)=502.8\] Now we compare this to\[K ^{2}=25/2025\]  So\[502.8\times2025=25\times x\]\[502.8\times2025\div25=x\]\[x= 40725.8\] but we will take it as \[x= 40725\]as the question demands it. Hope You understand.